Attachment 1 shows the case when excitation is sent at control point i.e. the PZT output. As before, free running laser noise in units of Hz/rtHz is added after the actuator and I've also shown shot noise being added just before the detector.
Again, we have a access to three output points for measurement. right at the output of mixer (the PDH error signal), the feedback signal to be applied by uPDH box (PZT Mon) and the output of the summing box SR560.
Doing loop algebra as before, we get:
So measurement of can be done by
This means at 100 Hz, with 10 integration cycles, we should have needed only 3 mV of excitation signal to get an SNR of 100. However, we have been unable to get good measurements with even 25 mV of excitation. We tried increasing the cycles, that did not work either.
This post is to summarize this analysis. We need more tests to get any conclusions.
I have added more degrees of freedom. The model includes x, y, z, pitch, yaw, roll and is controlled by a matrix of transfer functions (See Attachment 2). I have added 5 control filters to individually control UL, UR, LL, LR, and side. Eventually, this should become a matrix too but for the moment this is fine.
Note the Unit delay blocks in the control in Attachment 1. The model will not compile without these blocks.
Working in Xend with mask on has become unbearable. It is very hot there and I would really like if we fix this issue.
Today, the Xend Green laser was just unable to hold lock for longer than 10's of seconds. The longest I could see it hold lock was for about 2 minutes. I couldn't find anything obviously wrong with it. Attached are noise spectrums of error and control points. The control point spectrum shows good matching with typical free running laser noise.
Are the few peaks above 10 kHz in error point spectrum worrysome? I need to think more about it in a cooler place to make sure.
I wanted to take a high frequency spectrum of error point to make sure that higher harmonics of 250 kHz modulation frequency are not leaking into the PDH box after demodulation. However, the lock could not be maintained long enough to take this final measurement. I'll try again tomorrow morning. It is generally cooler in the mornings.
This post is just an update on what's happening. I need to work more to get some meaningful inferences about this loop.
I checked the BI situation on the HAM-A coil driver. It seems like these are sinking BIs and indeed need to be isolated from the Acromag unit GND to avoid contamination.
The BIs will have to be isolated on a different isolator. Now, the wires coming from the field (red) are connected to the second isolator's input and the outputs are connected to the Acromag BI module and the Acromag's RTN.
I updated the wiring diagram (attached) and the wiring spreadsheet.
In the diagram, you can notice that the BI isolator (the right one) is powered by the Acromag's +15V and switched when the coil driver's GND is supplied. I am not sure if it makes sense or not. In this configuration, there is a path between the coil driver's GND and the Acromag's GND but its resistance is at least 10KOhm. The extra careful option is to power the isolator by the coil driver's +V but there is no +V on any of the connectors going out of the coil driver.
I installed an additional isolator on the DIN rail and wired the remaining BOs (C1:SUS-ETMY_SD_ENABLE, C1:SUS-ETMY_LR_ENABLE) through it to the DB9F-4 feedthrough. I also added DB9F-3 for incoming wires from the RTS and made the required connection from it to DB9F-4.
I tested the new isolated BOs using the Windows machine (after stopping Modbus). As before, I measure the resistance between pin 5 (coil driver GND) and the channel under test. When I turn on the BO I see the resistance drops from inf to 166ohm and back to inf when I turn it off. Both channels passed the test.
Stephen and I discussed the in-vacuum OMC wiring.
- One of the OMCs has already been completed. (Blue)
- The other OMC is still being built. It means that these cables need to be built. (Pink)
- However, the cables for the former OMC should also be replaced because the cable harness needs to be replaced from the metal one to the PEEK one.
- The replacement of the harness can be done by releasing the Glenair Mighty Mouse connectors from the harness. (This probably requires a special tool)
- The link to the harness photo is here: https://photos.app.goo.gl/3XsUKaDePbxbmWdY7
- We want to combine the signals for the two OMCs into three DB25s. (Green)
- These cables are custom and need to be designed.
- The three standard aLIGO-style cables are going to be used. (Yellow)
- The cable stand here should be the aLIGO style.
Attachment 1 shows the closed loop of Xend Green laser Arm PDH lock loop. Free running laser noise gets injected at laser head after the PZT actuation as . The PDH error signal at output of miser is fed to a gain 1 SR560 used as summing junction here. Used in 'A-B mode', the B port is used for sending in excitation where .
We have access to three ports for measurement, marked at output of mixer, at output of SR560, and at PZT out monitor port in uPDH box. From loop algebra, we get following:
, where is the open loop transfer function of the loop.
So measurement of can be done in following two ways (not a complete set):
We did a quick check to make sure there is no saturation in the C1:SUS-ITMX_LSC_EXC analog path. For this, we looked at the SOS driver output monitors from the 1X4 chassis near MC2 on a scope. We found that even with 600 x 10 = 6000 counts of our 619 Hz excitation these outputs in particular are not saturating (highest mon signal was UL coil with 5.2 Vpp). In comparison, the calibration trials we have done before had 600 counts of amplitude, so we can safely increase our oscillator strength by that much
Things that remain to be investigated -->
I have attached an updated transfer function graph with the residual easier to see. I thought here I would include a better explanation of what this transfer function was measuring.
This transfer function was mainly about learning how to use DTT and Foton to make and measure transfer functions. Therefore it is just measuring across a single CDS filter block. X1SUP_ISOLATED_C1_SUS_SINGLE_PLANT_Plant_POS_Mod block to be specific. This measurement shows that the block is doing what I programmed it to do with Foton. The residual is probably just because the measured TF had fewer points than the calculated one.
The next step is to take a closed-loop TF of the system and the control module.
After that, I want to add more degrees of freedom to the model. both in the plant and in the controls.
We measured the Xend green laser PDH Open loop transfer function by following method:
I tested the digital inputs the following way: I connected a DB9 breakout to DB9M-5 and DB9M-6 where digital inputs are hosted. I shorted the channel under test to GND to turn it on.
I observed the channels turn from Disabled to Enabled using caget when I shorted the channel to GND and from Enabled to Disabled when I disconnected them.
I did this for all the digital inputs and they all passed the test.
I am still waiting for the other isolator to wire the rest of the digital outputs.
Next, I believe we should take some noise spectra of the Y end before we do the installation.
Tomorrow I will test the BIs using EPICs.
We attempted to simulate "oscillator based realtime calibration noise monitoring" in offline analysis with python. This helped us in finding about a factor of sqrt(2) that we were missing earlier in 16171. we measured C1:ALS-BEATX_FINE_PHASE_OUT_HZ_DQ when X-ARM was locked to main laser and Xend green laser was locked to XARM. An excitation signal of amplitude 600 was setn at 619 hz at C1:ITMX_LSC_EXC.
We got a value of . The calibration factor in use is from nm/cts from 13984.
Next steps could be to budget this noise while we setup some way of having this calibration factor generated in realitime using oscillators on a FE model. Calibrating actuation of a single optic in a single arm is easy, so this is a good test setup for getting a noise budget of this calibration method.
Added difference to the graph. I included the code so that others could see what it looks like and use it for easy use.
We found Mon7 in control room dead today afternoon. It's front power on green light is not lighting up. All other monitors are working as normal.
This monitor was used for looking at IMC camera analog feed. It is one of the most important monitors for us, so we should replace it with a different monitor.
Yehonathan and Paco disconnected the monitor and brought it down. We put it under the back table if anyone wants to fix it. Paco has ordered a BNC to VGA/HDMI converter to put in any normal monitor up there. It will happen this Wednesday. Meanwhile, I have changed the MON4 assignment from POP to Quad2 to be used for IMC.
I added a new XT1111 Acromag module to the c1auxey chassis. I sanitized and configured it according to the slow machines wiki instructions.
Since all the spare BIOs fit one DB37 connector I didn't add another feedthrough and combined them all on one and the same DB37 connector. This was possible because all the RTNs of the BIOs are tied to the chassis ground and therefore need only one connection. I changed the wiring spreadsheet accordingly.
I did a lot of rewirings and also cut short several long wires that were protruding from the chassis. I tested all the wires from the feedthroughs to the Acromag channels and fixed some wiring mistakes.
There is still an open issue with the BI channels not read by EPICS. They can still be read by the Windows machine though.
I looked into the issue that Yehonathan reported with the BI channels. I found the problem was with the .cmd file which sets up the Modbus interfacing of the Acromags to EPICS (/cvs/cds/caltech/target/c1auxey1/ETMYaux.cmd).
The problem is that all the channels on the XT1111 unit are being configured in Modbus as output channels. While it is possible to break up the address space of a single unit, so that some subset of channels are configured as inputs and another as outputs, I think this is likely to lead to mass confusion if the setup ever has to be modified. A simpler solution (and the convention we adopted for previous systems) is just to use separate Acromag units for BI and BO signals.
Accordingly, I updated the wiring plan to include the following changes:
So, one more Acromag XT1111 needs to be added to the c1auxey chassis, with the wiring changes as noted above. I have already updated the .cmd and EPICS database files in /cvs/cds/caltech/target/c1auxey1 to reflect these changes.
According to the BO interface circuit board https://dcc.ligo.org/D1001266, PCIN wires are connected to the coil driver and they are not pulled either way.
That means that they're either grounded or floating. I updated the drawing.
This RTS also use the BO interface with an opto isolator. https://dcc.ligo.org/LIGO-D1002593
Could you also include the pull up/pull down situations?
Since this Ocean Controls optoisolator has been shown to be compatible, I've gone ahead and ordered 10 more:
They are expected to arrive by Wednesday.
Here is an update and status report on the new BHD front-ends (FEs).
The changes to the FE BIOS settings documented in  do seem to have solved the timing issues. The RTS models ran for one week with no more timing failures. The IOP model on c1sus2 did die due to an unrelated "Channel hopping detected" error. This was traced back to a bug in the Simulink model, where two identical CDS parts were both mapped to ADC_0 instead of ADC_0/1. I made this correction and recompiled the model following the procedure in .
For lack of a better name, I had originally set up the user model on c1sus2 as "c1sus2.mdl" This week I standardized the name to follow the three-letter subsystem convention, as four letters lead to some inconsistency in the naming of the auto-generated MEDM screens. I renamed the model c1sus2.mdl -> c1su2.mdl. The updated table of models is below.
Renaming an RTS model requires several steps to fully propagate the change, so I've documented the procedure below for future reference.
On the target FE, first stop the model to be renamed:
controls@c1sus2$ rtcds stop c1sus2
Then, navigate to the build directory and run the uninstall and cleanup scripts:
controls@c1sus2$ cd /opt/rtcds/caltech/c1/rtbuild/release
controls@c1sus2$ make uninstall-c1sus2
controls@c1sus2$ make clean-c1sus2
Unfortunately, the uninstall script does not remove every vestige of the old model, so some manual cleanup is required. First, open the file /opt/rtcds/caltech/c1/target/gds/param/testpoint.par and manually delete the three-line entry corresponding to the old model:
If this is not removed, reinstallation of the renamed model will fail because its assigned DCUID will appear to already be in use. Next, find all relics of the old model using:
and manually delete each file and subdirectory containing the "sus2" name. Finally, rename, recompile, reinstall, and relaunch the model:
I used a tool developed by Chris, mdl2adl, to auto-generate a set of temporary sitemap/model MEDM screens. This package parses each Simulink file and generates an MEDM screen whose background is an .svg image of the Simulink model. Each object in the image is overlaid with a clickable button linked to the auto-generated RTS screens. An example of the screen for the C1BHD model is shown in Attachment 1. Having these screens will make the testing much faster and less user-error prone.
I generated these screens following the instructions in Chris' README. However, I ran this script on the c1sim machine, where all the dependencies including Matlab 2021 are already set up. I simply copied the target .mdl files to the root level of the mdl2adl repo, ran the script (./mdl2adl.sh c1x06 c1x07 c1bhd c1su2), and then copied the output to /opt/rtcds/caltech/c1/medm/medm_teststand. Then I redefined the "sitemap" environment variable on the chiara clone to point to this new location, so that they can be launched in the teststand via the usual "sitemap" command.
Currently, we are missing five 18-bit DACs needed to complete the c1sus2 system (the c1bhd system is complete). Since the first shipment, we have had no luck getting additional 18-bit DACs from the sites, and I don't know when more will become available. So, this week I took an inventory of all the 16-bit DACs available at the 40m. I located four 16-bit DACs, pictured in Attachment 2. Their operational states are unknown, but none were labeled as known not to work.
The original CDS design would call for 40 more 18-bit DAC channels. Between the four 16-bit DACs there are 64 channels, so if only 3/4 of these DACs work we would have enough AO channels. However, my search turned up zero additional 16-bit DAC adapter boards. We could check if first Rolf or Todd have any spares. If not, I think it would be relatively cheap and fast to have four new adapters fabricated.
DAQ network limitations and plan
To get deeper into the signal-integrity aspect of the testing, it is going to be critical to get the secondary DAQ network running in the teststand. Of all the CDS tools (Ndscope, Diaggui, DataViewer, StripTool), only StripTool can be used without a functioning NDS server (which, in turn, requires a functioning DAQ server). StripTool connects directly to the EPICS server run by the RTS process. As such, StripTool is useful for basic DC tests of the fast channels, but it can only access the downsampled monitor channels. Ian and Anchal are going to carry out some simple DAC-to-ADC loopback tests to the furthest extent possible using StripTool (using DC signals) and will document their findings separately.
We don't yet have a working DAQ network because we are still missing one piece of critical hardware: a 10G switch compatible with the older Myricom network cards. In the older RCG version 3.x used by the 40m, the DAQ code is hardwired to interface with a Myricom 10G PCIe card. I was able to locate a spare Myricom card, pictured in Attachment 3, in the old fb machine. Since it looks like it is going to take some time to get an old 10G switch from the sites, I went ahead and ordered one this week. I have not been able to find documentation on our particular Myricom card, so it might be compatible with the latest 10G switches but I just don't know. So instead I bought exactly the same older (discontinued) model as is used in the 40m DAQ network, the Netgear GSM7352S. This way we'll also have a spare. The unit I bought is in "like-new" condition and will unfortunately take about a week to arrive.
I mounted the optoisolator on the DIN rail and connected the 3 first channels
to the optoisolator inputs 1,3,4 respectively. I connected the +15V input voltage into the input(+) of the optoisolator.
The outputs were connected to DB9F-2 where those channels were connected before.
I added DB9F-1 to the front panel to accept channels from the RTS. I connected the fast channels to connectors 1,2,3 from DB9F-1 to DB9F-2 according to the wiring diagram. The GND from DB9F-1 was connected to both connector 5 of DB9F-2 and the output (-).
I tested the channels: I connected a DB9 breakout board to DB9F-2. I measured the resistance between the RTS GND and the isolated channels while switching them on and off. In the beginning, when I turned on the binary channels the resistance was behaving weird - oscillating between low resistance and open circuit. I pulled up the channels through a 100Kohm resistor to observe whether the voltage behavior is reasonable or not. Indeed I observed that in the LOW state the voltage between the isolated channel and slow GND is 15V and 0.03V in the HIGH state. Then I disconnected the pull up from the channels and measured the resistance again. It showed ~ stable 170ohm in the HIGH state and an open circuit in the LOW state. I was not able to reproduce the weird initial behavior. Maybe the optoisolator needs some warmup of some sort.
We still need to wire the rest of the fast channels to DBF9-3 and isolate the channels in DBF9-4. For that, we need another optoisolator.
I made a diagram (Attached). I think it explains the blue thing in the previous post.
I don't know what is the grounding situation in the RTS so I put a ground in both the coil driver and the RTS. Hopefully, only one of them is connected in reality.
- Could you explain what is the blue thing in Attachment 1?
- To check the validity of the signal chain, can you make a diagram summarizing the path from the fast BO - BO I/F - Acromag - This opto-isolator - the coil driver relay? (Cut-and-paste of the existing schematics is fine)
I borrowed the little red cart 🛒 to help clear the path for new optical tables in B252 West Bridge. Will return once I am done with it.
I fixed the PSL shutter button on Shutters summary page C1IOO_Mech_Shutter.adl. Now PSL switch changes C1:PSL-PSL_ShutterRqst channel. Earlier it was C1:AUX-PSL_ShutterRqst which doesn't do anything.
As Jon wrote we need to use the NPN configuration (see attachments). I tested the isolator channels in the following way:
1. I connected +15V from the power supply to the input(+) contact.
2. Signal wire from one of the digital outputs was connected to I1-4
3. When I set the digital output to HIGH, the LED on the isolator turns on.
4. I measure the resistance between O1-4 to output(-) and find it to be ~ 100ohm in the HIGH state and an open circuit in the LOW state, as expected from an open collector output.
Unlike the Acromag output, the isolator output is not pulled up in the LOW state. To do so we need to connect +15V to the output channel through a pull-up resistor. For now, I leave it with no pull-up. According to the schematics of the HAM-A Coil Driver, the digital output channels drive an electromagnetic relay (I think) so it might not need to be pulled up to switch back. I'm not sure. We will need to check the operation of these outputs at the installation.
During the testing of the isolator outputs pull-up, I accidentally ran a high current through O2, frying it dead. It is now permanently shorted to the + and - outputs rendering it unusable. In any case, we need another isolator since we have 5 channels we need to isolate.
I mounted the isolator on the DIN rail and started wiring the digital outputs into it. I connected the GND from the RTS to output(-) such that when the digital outputs are HIGH the channels in the coil driver will be sunk into the RTS GND and not the slow one avoiding GND contamination.
I was able to measure the transfer function of the plant filter module from the channel X1:SUP-C1_SUS_SINGLE_PLANT_Plant_POS_Mod_EXC to X1:SUP-C1_SUS_SINGLE_PLANT_Plant_POS_Mod_OUT. The resulting transfer function is shown below. I have also attached the raw data for making the graph.
Next, I will make a script that will make the photon filters for all the degrees of freedom and start working on the matrix version of the filter module so that there can be multiple degrees of freedom.
Following the conclusion, we are reverting the suspension gains to old values, i.e.
While the F2A filters, AC coil gains and input matrices are changed to as mentioned in 16066 and 16072.
The changes can be reverted all the way back to old settings (before Paco and I changed anything in the IMC suspensions) by running python scripts/SUS/general/20210602_NewIMCOldGains/restoreOldConfigIMC.py on allegra. The new settings can be uploaded back by running python scripts/SUS/general/20210602_NewIMCOldGains/uploadNewConfigIMC.py on allegra.
Unix Time = 1622676038
GPS Time = 1306711256
How about to use the non-Ag coated threaded shaft + the end SS masses with helicoils inserted? Does this save the masses to get stuck?
Can you just cut the viton tips smaller? If you cut it to have some wedge (or say, taper), it can get stuck with the vent hole.
Rana suggested in today's meeting to put in a notch filter in the XARM IR PDH loop to avoid suppressing the excitation line. We tried this today first with just one notch at 1069 Hz and then with an additional notch at 619 Hz and sent two simultaneous excitations.
I tried to push the clean Viton tips into the vented screws just to find out that the vented holes are too small. We need to drill 0.1" diameter holes about 0.1" deep into these screws and clean them again.
After changing the material of the Balance Mass from 6061 Al to 304 Steel, and changing the thickness to 0.21" from 0.25". The CoM is now 1.11mm below the clamping point.
Koji expected a mass change of ~ 4g to move the mass to 1.1mm. The 6061 mass weighed ~1.31g and the 304 mass weighs 4.1g.
A potential issue with this is the screw used the adjust the position of these balance masses, threads through both the aluminum ring and this now 304 steel mass. A non silver plated screw could cold weld at the mass, but a silver plated screw will gall in the aluminum threads.
The current vertical distance between the CoM and the wire clamping point on the 3" Ring assembly is 0.33mm. That is the CoM is .33 mm below the clamping point of the wire. I took the clamping point to be the top edge of the wire clamp piece. see the below attachments.
I am now modifying the dumbell mechanism at the bottom of the ring to move the CoM to the target distance of 1.1mm.
I attempted a single arm actuation calibration using IR beatnote (in the directions of soCal idea for DARM calibration)
An update on recent progress in the lab towards building and testing the new FEs.
The previously reported problem with the IOPs losing sync after a few minutes (16130) was resolved through a change in BIOS settings. However, there are many required settings and it is not trivial to get these right, so I document the procedure here for future reference.
The CDS group has a document (T1300430) listing the correct settings for each type of motherboard used in aLIGO. All of the machines received from LLO contain the oldest motherboards: the Supermicro X8DTU. Quoting from the document, the BIOS must be configured to enforce the following:
• Remove hyper-threading so the CPU doesn’t try to run stuff on the idle core, as hyperthreading simulate two cores for every physical core.
• Minimize any system interrupts from hardware, such as USB and Serial Ports, that might get through to the ‘idled’ core. This is needed on the older machines.
• Prevent the computer from reducing the clock speed on any cores to ‘save power’, etc. We need to have a constant clock speed on every ‘idled’ CPU core.
I generally followed the T1300430 instructions but found a few adjustments were necessary for diskless and deterministic operation, as noted below. The procedure for configuring the FE BIOS is as follows:
After completing the BIOS setup, I rebooted the new FEs about six times each to make sure the configuration was stable (i.e., would never hang during boot).
With the timing issue resolved, I proceeded to build basic user models for c1bhd and c1sus2 for testing purposes. Each one has a simple structure where M ADC inputs are routed through IIR filters to an M x N output matrix, which forms linear signal combinations that are routed to N DAC outputs. This is shown in Attachment 1 for the c1bhd case, where the signals from a single ADC are conditioned and routed to a single 18-bit DAC. The c1sus2 case is similar; however the Contec BO modules still needed to be added to this model.
The FEs are now running two models each: the IOP model and one user model. The assigned parameters of each model are documented below.
The user models were compiled and installed following the previously documented procedure (15979). As shown in Attachment 2, all the RTS processes are now working, with the exception of the DAQ server (for which we're still awaiting hardware). Note that these models currently exist only on the cloned copy of the /opt/rtcds disk running on the test stand. The plan is to copy these models to the main 40m disk later, once the new FEs are ready to be installed.
I installed several new AA and AI chassis in the test stand to interface with the ADC and DAC cards. This includes three 16-bit AA chassis, one 16-bit AI chassis, and one 18-bit AI chassis, as pictured in Attachment 3. All of the AA/AI chassis are powered by one of the new 15V DC power strips connected to a bench supply, which is housed underneath the computers as pictured in Attachment 4.
These chassis have not yet been tested, beyond verifying that the LEDs all illuminate to indicate that power is present.
I have received the opto-isolator needed to complete the new c1auxey system. I left it sitting on the electronics bench next to the Acromag chassis.
Here is the manufacturer's wiring manual. It should be wired to the +15V chassis power and to the common return from the coil driver, following the instructions herein for NPN-style signals. Note that there are two sets of DIP switches (one on the input side and one on the output side) for selecting the mode of operation. These should all be set to "NPN" mode.
Here's an updated X ARM ALS noise budget.
We went near the MC2 area and opened the lid to inspect the GigE and analog video monitors for MC2. Looked like whatever image is coming through the viewport is split into the GigE (for beam tracking) and the analog monitor. We hooked the monitor found on the floor nearby and tweaked the analog video camera around to get a feel for how the "ghost" image of the transmission moves around. It looks like in order to try and remove this "extra spots" we would need to tweak the beam tracking BS. We will consult the beam tracking authorities and return to this.
I was preparing a short write-up / test procedure for the custom HV coil driver, when I thought of something I can't resolve. I'm probably missing some really basic physics here - but why do we not account for the shot noise from DC current flowing through the series resistor? For a 4kohm resistor, the Johnson current noise is ~2pA/rtHz. This is the target we were trying to beat with our custom designed HV bias circuit. But if there is a 1 mA DC current flowing through this resistor, the shot noise of this current is 18pA/rtHz, which is ~9 times larger than the Johnson noise of the same resistor. One could question the applicability of this formula to calculate the shot noise of a DC current through a wire-wound resistor - e.g. maybe the electron transport is not really "ballistic", and so the assumption that the electrons transported through it are independent and non-interacting isn't valid. There are some modified formulae for the shot noise through a metal resistor, which evaluates to 10pA/rtHz for the same 4kohm resistor, which is still ~5x the Johnson noise.
In the case of the HV coil driver circuit, the passive filtering stage I added at the output to filter out the excess PA95 noise unwittingly helps us - the pole at ~0.7 Hz filters the shot noise (but not the Johnson noise) such that at ~10 Hz, the Johnson noise does indeed dominate the total contribution. So, for this circuit, I think we don't have to worry about some un-budgeted noise. However, I am concerned about the fast actuation path - we were all along assuming that this path would be dominated by the Johnson noise of the 4kohm series resistor. But if we need even 1mA of current to null some DC DARM drift, then we'd have the shot noise contribution become comparable, or even dominant?
I looked through the iLIGO literature, where single-stage suspensions were being used, e.g. Rana's manifesto, but I cannot find any mention of shot noise due to DC current, so probably there is a simple explanation why - but it eludes me, at least for the moment. The iLIGO coil drivers did not have a passive filter at the output of the coil driver circuit (at least, not till this work), and there isn't any feedback gain for the DARM loop at >100 Hz (where we hope to measure squeezing) to significantly squash this noise.
Attachment #1 shows schematic topologies of the iLIGO and proposed 40m configs. It may be that I have completely misunderstood the iLIGO config and what I've drawn there is wrong. Since we are mainly interested in the noise from the resistor, I've assumed everything upstream of the final op-amp is noiseless (equivalently, we assume we can sufficiently pre-filter these noises).
Attachment #2 shows the relative magnitudes of shot noise due to a DC current, and thermal noise of the series resistor, as a function of frequency, for a few representative currents, for the slow bias path assuming a 0.7Hz corner from the 4kohm/3uF RC filter at the output of the PA95.
Some lit review suggests that it's actually pretty hard to measure shot noise in a resistor - so I'm guessing that's what it is, the mean free path of electrons is short compared to the length of the resistor such that the assumption that electrons arrive independently and randomly isn't valid. So Ohm's law dictates and that's what sets the current noise. See, for example, pg 432 of Horowitz and Hill.
Here is our first attempt at a single-arm noise budget for ALS.
Attachment 1 shows the loop diagram we used to calculate the contribution of different noises.
Attachment 2 shows the measured noise at C1:ALS-BEATX_PHASE_FINE_OUT_HZ when XARM was locked to the main laser and Xend Green laser was locked to XARM.
All remaining chasses have been reworked and placed on the floor along the west wall in Room 104.
The test was succesful and brought back the IMC to lock point at the end.
We calculated new input matrix using same code in scripts/SUS/InMatCalc/sus_diagonalization.py . Attachment 1 shows the results.
The calculations are present in scripts/SUS/InMatCalc/MC1.
We uploaded the new MC1 input matrix at:
Unix Time = 1621963200
GPS Time = 1305998418
This was done by running python scripts/SUS/general/20210525_NewMC1Settings/uploadNewConfigIMC.py on allegra. Old IMC settings (before Paco and I started workin on 40m) can be restored by running python scripts/SUS/general/20210525_NewMC1Settings/restoreOldConfigIMC.py on allegra.
Everything looks as stable as before. We'll look into long term trends in a week to see if this helped at all.
Differential misalignment of the OMCs
40m BHD will employ two OMCs on the BHD platform. We will have two SOSs for each of the LO and AS beams. The challenge here is that the input beam must optimally couple to the OMCs simultaneously. This is not easy as we won't have independent actuators for each OMC. e.g. The alignment of the LO beam can be optimally adjusted to the OMC1, but this, in general, does not mean the beam is optimally aligned to the OMC2.
When a beam with the matched mode to an optical cavity has a misalignment, the power coupling C can be reduced from the unity as
where is the waist radius, is the divergence angle defined as , and are the beam lateral translation and rotation at the waist position.
The waist size of the OMC is 500um. Therefore = 500um and = 0.68 mrad. If we require C to be better than 0.995 according to the design requirement document (T1900761). This corresponds to (only) to be 35um and (only) to be 48urad. These numbers are quite tough to be realized without post-installation adjustment. Moreover, the OMCs themselves have individual differences in the beam axis. So no matter how we set the mechanical precision of the OMC installation, we will introduce a maximum of 1mm and ~5mrad uncertainty of the optical axis.
Suppose we adjust the incident beam to the OMC placed at the transmission side of the BHD BS. The reflected beam at the BS can be steered by picomotors. The distance from the BS to the OMC waist is 12.7" (322mm) according to the drawing.
So we can absorb the misalignment mode of (, ) = (0.322 , ). This is a bit unfortunate. 0.322m is about 1/2 of the rayleigh range. Therefore, this actuation is still angle-dominated but a bit of translation is still coupled.
If we enable to use the third picomotor on the BHD BS mount, we can introduce the translation of the beam in the horiz direction. This is not too huge therefore we still want to prepare the method to align the OMC in the horiz direction.
The difficult problem is the vertical alignment. This requires the vertical displacement of the OMC. And we will not have the option to lower the OMC. Therefore if the OMC2 is too high, we have to raise the OMC1 so that the resulting beam is aligned to the OMC2. i.e. we need to maintain the method to raise both OMCs. (... or swap the OMCs). From the images of the OMC beam spots, we'll probably be able to analyze the intracavity axes of the OMCs. So we can always place the OMC with a higher optical axis at the transmission side of the BHD BS.
We've set a free swing test to trigger at 3:30 am tomorrow for MC1. The script for tests is running on tmux session named 'freeSwingMC1' on rossa. The script will run for about 4.5 hrs and we'll correct the input matrix tomorrow from the results. If anyone wants to work during this time (3:30 am to 8:00 am), you can just kill the script by killing tmux session on rossa. ssh into rossa and type tmux kill-session -t freeSwingMC1.
We should redo the MC1 input matrix optimization and the coil balancing afterward as we did everything based on the noisy UL OSEM values.
Updated IOO.strip on Zita to show WFS2 pitch and yaw trends (C1:IOO-WFS2_PIY_OUT16 and C1:IOO-WFS2_YAW_OUT16) and changed the colors slightly to have all pitch trends in the yellow/brown band and all yaw trends in the pink/purple band.
No one says, "Here I am attaching a cool screenshot, becuz else where's the proof? Am I right or am I right?"
Mon May 24 18:10:07 2021 [Update]
After waiting for some traces to fill the screen, here is a cool screenshot (Attachment 1). At around 2:30 PM the MC unlocked, and the BS_Z (vertical) seismometer readout jumped. It has stayed like this for the whole afternoon... The MC eventually caught its lock and we even locked XARM without any issue, but something happened in the 10-30 Hz band. We will keep an eye on it during the evening...
Tue May 25 08:45:33 2021 [Update]
At approximately 02:30 UTC (so 07:30 PM yesterday) the 10-30 Hz seismic step dropped back... It lasted 5 hours, mostly causing BS motion along Z (vertical) as seen by the minute trend data in Attachment 2. Could the MM library have been shaking? Was the IFO snoring during its afternoon nap?
- High priority units: 2x 18AI / 1x 16AI / 3x 16AA
All six are reworked and on the electronics workbench. The rest should be ready by the end of the week.
The new HAM-A coil drivers have a single DB9 connector for all the binary inputs. This requires that the dewhitening switching signals from the fast system be spliced with the coil enable signals from c1auxey. There is a common return for all the binary inputs. To avoid directly connecting the grounds of the two systems, I have looked for a suitable opto-isolator for the c1auxey signals.
I best option I found is the Ocean Controls KTD-258, a 4-channel, DIN-rail-mounted opto-isolator supporting input/output voltages of up to 30 V DC. It is an active device and can be powered using the same 15 V supply as is currently powering both the Acromags and excitation. I ordered one unit to be trialed in c1auxey. If this is found to be good solution, we will order more for the upgrades of c1auxex and c1susaux, as required for compatibility with the new suspension electronics.
The plant transfer function of the pendulum in the s domain is:
Using Foton to make a plot of the TF needed and using m=40kg, w0=3Hz, and Q=50 (See attachment 1). It is easiest to enter the above filter using RPoly and saved it as Plant_V1